Suppose you are performing a one-way ANOVA with only the information given in the following table.

Sources of Variation

Degrees of Freedom

Sum of Squares

Between

4

200

Within

45

3547

Thep-value for the test of the equality of the means of all populations is calculated to be0.6406. Suppose you plan to increase the sample sizes for all groups but keep them all equal. However, when you do this, the sum of squares within samples and the sum of squares between samples (magically) remain the same. What are the smallest sample sizes for groups that would make this result significant at the5%significance level?

Sample sizes=Enter your answer in accordance to the question statement
Answer #1

solution

total df=49

number of groups =4+1=5

number of points = df+1

=49+1

=50

group sample size=no of points /number ofgroups

=50/5

=10